Diego Company manufactures one product that is sold for $75 per unit in two geographic regions—the East and West regions. The following information pertains to the company’s first year of operations in which it produced 46,000 units and sold 42,000 units.

Variable costs per unit:
Manufacturing:
Direct materials $ 25
Direct labor $ 20
Variable manufacturing overhead $ 2
Variable selling and administrative $ 4
Fixed costs per year:
Fixed manufacturing overhead $ 644,000
Fixed selling and administrative expense $ 388,000

The company sold 31,000 units in the East region and 11,000 units in the West region. It determined that $200,000 of its fixed selling and administrative expense is traceable to the West region, $150,000 is traceable to the East region, and the remaining $38,000 is a common fixed expense. The company will continue to incur the total amount of its fixed manufacturing overhead costs as long as it continues to produce any amount of its only product.

Foundational 4-14 (Algo)

14a. Diego is considering eliminating the West region because an internally generated report suggests the region’s total gross margin in the first year of operations was $46,000 less than its traceable fixed selling and administrative expenses. Diego believes that if it drops the West region, the East region's sales will grow by 5% in Year 2. Using the contribution approach for analyzing segment profitability and assuming all else remains constant in Year 2, what would be the profit impact of dropping the West region in Year 2?

15b, Assume the West region invests $36,000 in a new advertising campaign in Year 2 that increases its unit sales by 20%. If all else remains constant, what would be the profit impact of pursuing the advertising campaign?
Answer #1
Selling price per unit 75
Less: Variable expense
Direct materials 25
Direct labor 20
Variable manufacturing overhead 2
Variable selling and administrative 4
Total Variable expense 51
Contribution margin per unit 24
14a
Loss in Contribution margin of West -264000 =11000*24
Increase in Contribution margin of East 37200 =31000*5%*24
Avoidable fixed selling and administrative expense 200000
Net change in Profits -26800
Profits will decrease by 26800
15b
Increase in unit sales 2200 =11000*20%
X Contribution margin per unit 24
Increase in Contribution margin 52800
Less: Advertising costs -36000
Net change in Profits 16800
Profits will increase by 16800